OpenStudy (anonymous):

What is 3/4 + N = 19/20

OpenStudy (anonymous):

I would first make it the same denominator so 3/4 could be multiplied by 5 to be 15/20 15/20 + n = 19/20 Subtract 15/20 from both sides -15/20 -15/20 --------------------- n = 4/20

OpenStudy (anonymous):

(3/4)20/20+N=(19/20)4/4 60/80+N=76/80 N=16/80=8/40=2/10=1/5

OpenStudy (anonymous):

IF you want it simplified it would be 1/5